OpenStudy (anonymous):

The sum of two consecutive odd integers is 24. Find the integers. Let n = first integer. (The second let statement is not given. You will have to decide what it should be.)

OpenStudy (anonymous):

Which equation should be used to solve this problem?

OpenStudy (merchandize):

11&13

OpenStudy (merchandize):

11+13=24

OpenStudy (lalaly):

n+(n+2)=24

OpenStudy (merchandize):

wo consecutive odd integers Let x and (x+2)represent the two consecutive odd integers Question states*** x +(x+2) = 24 Solving for x 2x = 22 x = 11 The two consecutive odd integers are 11,13 CHECKING our Answer*** 11 + 13 = 24

OpenStudy (anonymous):

^Thats what I was looking 4 thanks :)

OpenStudy (lalaly):

:)

OpenStudy (anonymous):

@lalaly :) Thanks

OpenStudy (lalaly):

you're welcome:D